# Physics-Informed Recurrent Network for Gas Pipeline Network Parameters Identification

### v1

@@ Abstract
As a part of the integrated energy system (IES), gas pipeline networks can provide additional flexibility to power systems through coordinated optimal dispatch. An accurate pipeline network model is critical for the optimal operation and control of IESs. However, inaccuracies or unavailability of accurate pipeline parameters often introduce errors in the mathematical models of such networks. This paper proposes a physics-informed recurrent network (PIRN) model to identify the state-space model of gas pipelines. The approach combines data-driven learning from measurement data with the fluid dynamics described by partial differential equations. By embedding the physical state-space model within the recurrent network, parameter identification is transformed into a training process for a PIRN. Similar to standard recurrent neural networks, this model can be implemented using the PyTorch framework and trained via backpropagation. Case studies demonstrate that our method accurately estimates gas pipeline models from sparse terminal node measurements, providing robust performance and significantly higher parameter efficiency. Furthermore, the identified models can be seamlessly integrated into optimization frameworks.
keywords:
@@
Integrated energy system , Physics-informed , Recurrent network , Gas pipeline network , Parameter identification
journal: Energy
[…]
[thu]organization=State Key Laboratory of Power Systems, Department of Electrical Engineering, addressline=Tsinghua University, postcode=100084, state=Beijing, country=China
[…]
[hkp]organization=Hong Kong Polytechnic University, state=Hong Kong Special Administrative Region, country=China
[…]
[byte]organization=ByteDance Ltd., postcode=100083, state=Beijing, country=China
{highlights}
A physics-informed recurrent network (PIRN) is proposed, integrating fluid dynamics with data-driven learning.
The PIRN model requires only sparse terminal node measurements to identify the state-space parameters of gas pipelines.
The PIRN model demonstrates superior parameter efficiency and can be seamlessly integrated into optimization frameworks.
1 Introduction
1.1 Background and Motivation
@@
In recent years, significant efforts have been directed toward advancing energy conversion technologies within integrated energy systems [1], with the dual goals of improving energy efficiency and reducing overall carbon emissions [2]. Gas-fired units and power-to-gas devices are key examples of these technologies and have seen widespread deployment [3]. To ensure economically optimal operation, the coordinated dispatch of gas and power systems is essential [4]. However, inaccuracies in the mathematical modeling of gas pipeline networks can render the results of gas system optimal dispatch calculations infeasible.
@@
In the context of optimal dispatch, it is critical to account for the dynamic behavior of gas flow [5], which is governed by the partial differential equations (PDEs) of fluid dynamics [6]. The parameters that define this dynamic process are directly influenced by the gas pipeline network’s characteristics. A promising alternative to deriving these parameters analytically is the use of data-driven approaches [7].
@@
Recurrent Neural Networks (RNNs) are a commonly used type of artificial neural network, particularly well-suited for processing sequential and time-dependent data [8]. Their architecture includes directed cycles, which enable RNNs to maintain a hidden state, effectively capturing and retaining information from previous inputs. This memory capability makes RNNs highly appropriate for simulating dynamic, state-dependent systems, such as gas pipeline networks. However, the inherent ”black box” nature of traditional neural networks presents challenges in terms of interpretability, making it difficult to integrate them seamlessly into optimization problems, such as those encountered in optimal dispatch.
@@
In recent years, physics-informed neural networks (PINNs) [9] have garnered significant attention within the research community. A PINN seamlessly integrates domain knowledge, particularly physics-based equations, into the standard neural network framework, making it a powerful tool for the simulation, parameter identification, and control of complex physical systems [10]. This physics knowledge can inform the design of loss functions, initialization of model parameters, and the architecture of the neural network [11]. The inspiration from PINNs provide possible solutions to leverage both the data-driven method and the physics model to identify the parameters of gas pipeline network and help it integrated in the optimization problem.
1.2 Literature Review and Research Gap
@@
Optimization-based approaches aim to identify parameters by minimizing an objective function that quantifies error. For instance, a reduced-order model was developed in [12] and [13], where nonlinear programming and a least squares objective function were employed to estimate both the dynamic states and parameters of gas pipeline networks. Similarly, maximum likelihood and least squares methods were utilized to identify parameters in [14]. A Bayesian approach to parameter estimation, specifically targeting the friction coefficient in gas networks, was proposed in [15]. Additionally, [16] employed a lumped-parameter model under both stationary and non-stationary gas flows to estimate pipeline parameters. The work in [17] presented an optimization problem to determine the steady-state of gas networks, where a neural network was used to initialize the optimization process and expedite its solution.
@@
Kalman filtering is another effective method, particularly for dealing with noisy measurements in parameter estimation. For example, [21, 22] utilized the Kalman filter to estimate the dynamic states of natural gas pipelines. Furthermore, a robust Kalman filter-based approach, designed to mitigate the impact of poor-quality data, was introduced in [23]. The work in [24] proposed a graph-frequency domain Kalman filter that reduces the influence of outliers in the data.
@@
In summary, much of the existing research focuses on simulating the dynamic behavior of gas networks, often incorporating physical information to enhance model performance [27, 28]. However, less attention has been directed towards parameter identification in gas networks—a process that is essentially the inverse of simulation. Furthermore, current data-driven solutions for gas pipeline parameter estimation tend to focus on steady-state conditions, with limited attention paid to dynamic state transitions, which are either oversimplified or ignored. This shortcoming restricts their applicability to real-world gas system operations. For example, the widely used quasi-dynamic models, which rely on terminal node states to represent the entire pipeline, can introduce significant inaccuracies in optimal dispatch problems [29]. Additionally, traditional neural network models face challenges in safety-critical applications due to their poor interpretability. Thus, developing interpretable models capable of accurately capturing the physical dynamics of gas pipeline networks is critical for improving operational efficiency and safety.
1.3 Method and Contributions
To identify the state-space model of gas pipeline networks and integrate the results into optimization problems, this paper introduces a physics-informed recurrent network model. Unlike standard RNNs that rely on artificial neural perceptron units, our model incorporates a physical state-space representation of the gas pipeline network. This transforms the black-box perceptron units in traditional RNNs into a physically interpretable gray-box model, where the parameters correspond directly to those in the dynamic equation matrices, providing a higher degree of interpretability. In this framework, the task of identifying pipeline parameters is reformulated as a training process for the physics-informed recurrent network. Leveraging the memory capability of RNN architecture, our method requires only sparse measurements from the terminal nodes of the gas pipeline, while the internal states along the pipeline are estimated and stored in the hidden layers of the network. Embedding prior physical knowledge prevents the network from converging to physically unrealistic local optima, ensuring both stability and physical plausibility. Consequently, the identified state-space model can be seamlessly integrated into any optimal dispatch problem for gas networks.
@@ 1.3 Method and Contributions
(1) We propose a novel physics-informed recurrent network method for identifying the state-space model of gas pipeline networks. Unlike standard RNN models, where perceptron units are used, we substitute these units with the physical state-space functions of the gas pipeline network, transforming the parameter identification process into a training task.
(2) The proposed method relies solely on sparse measurement data from terminal nodes of each pipeline, while the internal gas states along the pipeline are estimated and stored in the hidden layers of the physics-informed recurrent network.
(3) Our physics-informed recurrent network model can be efficiently implemented in the PyTorch framework with minimal modifications to the standard RNN architecture. Compared to traditional RNNs, this approach offers improved robustness and significantly higher parameter efficiency.
The remainder of this paper is organized as follows. Section 2 provides an overview of the problem formulation and methodology. The detailed state-space model identification method is presented in Section 3. Numerical tests are performed in Section 4 and conclusions are made in Section 5.
2 Problem Formation and Methodology Overview
The diagram of the gas pipeline network and its measurement configuration are illustrated in Figure 1. Natural gas originates from gas wells (GWs) and power-to-gas devices (P2Gs). Pressurized by compressors, natural gas is transported to end-users, such as gas turbines (GTs), via the pipeline network. The transportation of gas through the pipelines is a relative slow dynamic process [30]. The gas pressure and mass flow rate at a specific time and position are two physical quantities that characterize the dynamic states of gas within the pipeline network. These states are governed by the following fluid partial differential equations:
@@ 2 Problem Formation and Methodology Overview
To establish the state-space model for the entire gas pipeline network, we discretize the spatial locations in the partial differential equations (1)-(2). As shown in Figure 1, to depict gas states along a pipeline, it is partitioned into equidistant discrete computation nodes. The states at internal computation nodes are unobservable; only the states at terminal nodes can be measured. Thus, we can only leverage the measurement data at terminal nodes to identify the state-space model of the entire pipeline network. Additionally, the dynamic process of gas cannot be expressed merely as a state transition function due to topological constraints inherent in the gas network, which must be maintained throughout the dynamic process. These characteristics present significant technical challenges in identifying the state-space model of the gas pipeline network.
Leveraging the memory retention feature of RNNs, this paper presents a physics-informed recurrent network model to identify the state-space model of gas pipeline network with sparse measurement data. The embedding of physics state-space model inside the recurrent network model make it physically interpretable and can be integrated into arbitrary optimal dispatch problems. A comparison between standard RNN and physics-informed recurrent network model is shown in Table 1.
To demonstrate the effectiveness of the identified state-space model, the optimal dispatch of the integrated power-gas system is implemented. By jointly optimizing the operation of generators, GTs, GWs, P2Gs and compressors, the economic benefits of the integrated system can be fully exploited.
Table 1: Comparison between Standard RNN and Physics-Informed Recurrent Network
Comparison items Standard RNN Physics-informed recurrent network
Output layer Predict data Predict measurable states of gas
Label Ground truth Measurement data of measurable variables
Training algorithm Backward propagation through time (BPTT)
Implementation framework PyTorch/TensorFlow
3 State-Space Model Identification of Gas Pipeline Network
3.1 Mathematical Model of Gas Pipeline Network
To derive a more practical state-space model of a gas pipeline network, simplify and discrete the fluid partial differential equations (1)-(2) and supplement the topology constraints of the whole network.
@@ 3.1 Mathematical Model of Gas Pipeline Network
where $u$ denotes the speed of sound; $A$ denotes the cross-section area of pipeline. Equation (6) holds under approximately adiabatic condition.
Finally, based on the approximation in (3)-(5) and the variables replacement in (6)-(7), the partial differential equations (1)-(2) are transformed into
@@
$\left(1+\frac{\lambda_{l}v_{b,l}\Delta t}{D_{l}}\right)f_{i,t}-\frac{A_{l}\Delta t}{\Delta x_{l}}\pi_{i-1,t}+\left(\frac{A_{l}\Delta t}{\Delta x_{l}}-\frac{\lambda_{l}A_{l}v_{b,l}^{2}\Delta t}{2u^{2}D_{l}}\right)\pi_{i,t}=f_{i,t-1}$ (11)
where $\mathcal{L}$ and $\mathcal{S}$ denote the index set of pipelines and computation nodes, respectively; $\mathcal{S}_{l}$ denotes the index set of computation node in the pipeline $l$. For the convenience of numerical computation, each pipeline is divided to sections by several computation nodes.
@@
3.2 State-space Model of Gas Pipeline Network
@@
To demonstrate the construction of state-space model, we use a 3-node small gas system as an example, as shown in Figure 2. Each pipeline is divided into 2 segments and there are total of 6 computation points, numbered $1^{\prime}$-$6^{\prime}$ , respectively.
@@
The dynamic function (18) of this 3-node simple gas system can be written in the following matrix form as shown in Figure 3.
@@
The detail parameters inside the sparse matrix are as follows:
@@
According to the matrix dynamic function in Figure 4, to calculate the state variables $\bm{h}_{t}$, it only needs the values in $\bm{h}_{t-1}$ that start from the second computation points. This is because the state of the first computation points of each pipeline ($1^{\prime}$ and $4^{\prime}$ in this case) are not independent. Their values are also determined by topology constraints at terminal nodes. The constant matrix $\bm{S}$ is used to select the independent states from all state variables. In this small case, the value of matrices $\bm{S}$ and $\bm{H}$ are shown in Figure 4.
3.3 Identify the State-Space Model with Physics-Informed Recurrent Network
This subsection introduces the state-space model of the pipelines using the measurement data in the dynamic process. A novel physics-informed recurrent network is proposed to identify the parameters in the state-space model (18)-(19).
@@ 3.3 Identify the State-Space Model with Physics-Informed Recurrent Network
Base on the state transition equation (26) and the output equation (19), the diagram of physics-informed recurrent network model training architecture is shown in Figure 5. The control variables $\bm{u}_{t}$ are inputs of this system. The state variables $\bm{h}_{t}$ updates based on the input variables and the previous state $\bm{h}_{t-1}$. The outputs $\hat{\bm{y}}_{t}$ are the predicted output data and the labels $\bm{y}_{t}$ are the real measured states.
Figure 5: Diagram of physics-informed recurrent network framework.
@@
where $\alpha_{\theta}$ and $\alpha_{h}$ denote the learning rates of $\hat{\bm{\theta}}$ and $\bm{h}_{0}$, respectively, and they are the hyper-parameters in the model. With the PyTorch framework, we leverage automatic differentiation to compute partial derivatives numerically, facilitating backward propagation and parameter updates in (29)-(30). However, this process involves repeated inversion of the large-scale sparse matrix $\bm{K}(\hat{\bm{\theta}})$ to obtain updated $\bm{J}$ as the parameters $\hat{\bm{\theta}}$ updated, which introduces a significant computational burden and can lead to numerical instability. To address this issue, we apply the matrix inversion lemma to update the new value of the matrix $\bm{J}$. When the parameters $\hat{\bm{\theta}}$ inside the sparse matrix $\bm{K}(\hat{\bm{\theta}})$ change to $\hat{\bm{\theta}}+\Delta\hat{\bm{\theta}}$, the matrix update can be denoted as:
@@
3.4 Training Techniques of Physics-Informed Recurrent Network
(1) Warm start
A warm start avoids numerical instability during matrix inversion more efficiently than random initialization, as well as accelerates the training process. Since all the parameters have an explicit physical meaning, the initial values of pending parameters $\hat{\bm{\theta}}$ and initial states $\bm{h}_{0}$ can be estimated based on the prior knowledge of physical meanings. The initial value of pending parameters $\hat{\bm{\theta}}$ can be directly set as $\bm{\theta}_{0}$. The initial states $\bm{h}_{0}$ in steady state can be estimated according to the physical model of gas pipeline. According to the partial differential equations of gas pipelines, the expressions of mass flow and gas pressure distribution along the pipeline in steady state are as follows:

### v2

# Physics-Informed Recurrent Network for State-Space Modeling of Gas Pipeline Networks

@@ Abstract
As a part of the integrated energy system (IES), gas pipeline networks can provide additional flexibility to power systems through coordinated optimal dispatch. An accurate pipeline network model is critical for the optimal operation and control of IESs. However, inaccuracies or unavailability of accurate pipeline parameters often introduce errors in the state-space models of such networks. This paper proposes a physics-informed recurrent network (PIRN) to identify the state-space model of gas pipelines. It fuses sparse measurement data with fluid-dynamic behavior expressed by partial differential equations. By embedding the physical state-space model within the recurrent network, parameter identification becomes an end-to-end PIRN training task. The model can be realized in PyTorch through modifications to a standard RNN backbone. Case studies demonstrate that our proposed PIRN can accurately estimate gas pipeline models from sparse terminal node measurements, providing robust performance and significantly higher parameter efficiency. Furthermore, the identified state-space model of the pipeline network can be seamlessly integrated into optimization frameworks.
Index Terms:
@@
Gas pipeline network, Integrated energy system, Parameter identification, Physics-informed, Recurrent network
I Introduction
I-A Background and Motivation
@@
Advancements in integrated energy systems have focused on optimizing energy conversion to enhance efficiency and lower carbon emissions. Gas-fired units and power-to-gas devices are key examples of these technologies and have seen widespread deployment. To ensure economically optimal operation, the coordinated dispatch of gas and power systems is essential. However, inaccuracies in the mathematical modeling of gas pipeline networks can render the results of gas system optimal dispatch calculations infeasible.
@@
In the context of optimal dispatch, it is critical to account for the dynamic behavior of gas flow [1], which is governed by the partial differential equations (PDEs) of fluid dynamics. The parameters that define this dynamic process are directly influenced by the gas pipeline network’s characteristics. A promising alternative to deriving these parameters analytically is the use of data-driven approaches [2].
@@
Recurrent Neural Networks (RNNs) are a commonly used type of artificial neural network, particularly well-suited for processing sequential and time-dependent data [3]. Their architecture includes directed cycles, which enable RNNs to maintain a hidden state, effectively capturing and retaining information from previous inputs. This memory capability makes RNNs highly appropriate for simulating dynamic, state-dependent systems, such as gas pipeline networks. However, the inherent black-box nature of traditional neural networks presents challenges in terms of interpretability, making it difficult to integrate them seamlessly into optimization problems, such as those encountered in optimal dispatch.
@@
In recent years, physics-informed neural networks (PINNs) [4] have garnered significant attention within the research community. A PINN can seamlessly integrate domain knowledge, particularly physics-based equations, into the classic neural network framework, making it a powerful tool for the simulation, parameter identification, and control of complex physical systems [5]. This physics knowledge can inform the design of loss functions, initialization of model parameters, and the architecture of the neural network [6]. The inspiration from PINNs provides possible solutions to leverage both the data-driven method and the physics model to identify the parameters of a gas pipeline network.
This paper introduces a physics-informed recurrent network (PIRN) for learning the state-space model of gas pipeline networks. By embedding the physical state-space dynamics into each recurrent unit, PIRN unites gray-box interpretability with the capacity to infer full internal states from sparse terminal measurements. Its streamlined architecture and transparent structure further enable the identified model to integrate seamlessly into downstream optimization tasks.
I-B Literature Review and Research Gap
@@
Optimization-based approaches aim to identify parameters by minimizing an objective function that quantifies error. For instance, a reduced-order model was developed in [7], where nonlinear programming and least squares objective function were employed to estimate both the dynamic states and parameters of gas pipeline networks. Similarly, maximum likelihood and least squares methods were utilized to identify parameters in [8]. A Bayesian approach to parameter estimation, specifically targeting the friction coefficient in gas networks, was proposed in [9]. Additionally, [10] employed a lumped-parameter model under both stationary and non-stationary gas flows to estimate pipeline parameters. The work in [11] presented an optimization problem to determine the steady-state of gas networks, where a neural network was used to initialize the optimization process and expedite its solution.
@@
Kalman filtering is another effective method, particularly for dealing with noisy measurements in parameter estimation. For example, [15] utilized the Kalman filter to estimate the dynamic states of natural gas pipelines. Furthermore, a robust Kalman filter-based approach, designed to mitigate the impact of poor-quality data, was introduced in [chenRobustKalmanFilterBased2021]. The work in [16] proposed a graph-frequency domain Kalman filter that reduces the influence of outliers in the data.
@@
In summary, much of the existing research concentrates on simulating the dynamic behavior of gas networks, often integrating physical information to improve model performance [19]. In contrast, relatively little attention has been devoted to parameter identification in gas pipeline networks, a task that fundamentally involves reversing the simulation process. Moreover, current data-driven approaches to parameter estimation typically emphasize steady-state conditions, offering limited treatment of dynamic state transitions, which are often oversimplified or entirely neglected. As a result, these methods face significant limitations when applied to real-world gas system operations. For example, the widely used quasi-dynamic models, which rely on terminal node states to represent the entire pipeline, can introduce significant inaccuracies in optimal dispatch problems [20]. Additionally, traditional neural network models face challenges in safety-critical applications due to their poor interpretability. Thus, developing interpretable models capable of accurately capturing the physical dynamics of gas pipeline networks is critical for improving operational efficiency and safety.
I-C Methods and Contributions
To identify the state-space model of gas pipeline networks and integrate the results into optimization problems, this paper introduces a PIRN model. Unlike classic RNNs that rely on artificial neural perceptron units, our model incorporates a physical state-space representation of the gas pipeline network. This transforms the black-box perceptron units in classic RNNs into a physically interpretable gray-box model, where the parameters correspond directly to those in the dynamic equation matrices, providing a higher degree of interpretability. In this framework, the task of identifying pipeline parameters is reformulated as a training process for the PIRN. Leveraging the memory capability of RNN architecture, our method requires only sparse measurements from the terminal nodes of the gas pipeline, while the internal states along the pipeline are estimated and stored in the hidden layers of the network. Embedding prior physical knowledge prevents the network from converging to physically unrealistic local optima, ensuring both stability and physical interpretability. Consequently, the identified state-space model can be seamlessly integrated into any optimal dispatch problem for gas networks.
@@ I-C Methods and Contributions
(1) We introduce a novel physics-informed recurrent network (PIRN) architecture for learning the state-space dynamics of gas pipeline networks. Rather than using generic perceptron units, our approach replaces them with physical state-space functions derived from the governing gas-flow equations, casting model identification as a structured, physics-grounded learning task.
(2) The proposed PIRN requires only sparse measurements at pipeline terminal nodes. Internal gas states are inferred and encoded within the network’s hidden layers, enabling full-state reconstruction from highly sparse observations.
(3) Our model integrates seamlessly with the PyTorch framework, requiring only minor adaptations to classic RNN structures. Compared to the classic RNNs, it achieves superior robustness, enhanced interpretability, and significantly improved parameter efficiency.
The remainder of this paper is organized as follows. Section II provides an overview of the problem formulation and methodology. The detailed state-space model identification method is presented in Section III. Numerical tests are performed in Section IV and conclusions are made in Section V.
II Problem Formation
The diagram of the gas pipeline network and its measurement configuration are illustrated in Fig. 1. Natural gas originates from gas wells (GWs) and power-to-gas devices (P2Gs). Pressurized by compressors, natural gas is transported to end-users, such as gas turbines (GTs), via the pipeline network. The transportation of gas through the pipelines is a relatively slow dynamic process [21]. The gas pressure and mass flow rate at a specific time and position are two physical quantities that characterize the dynamic states of gas within the pipeline network. These states are governed by the following fluid partial differential equations:
@@ II Problem Formation
To establish the state-space model for the entire gas pipeline network, we discretize the spatial locations in the partial differential equations (1)-(2). As shown in Fig. 1, to depict gas states along a pipeline, it is partitioned into equidistant discrete computation nodes. The states at internal computation nodes are unobservable; only the states at terminal nodes can be measured. Thus, we can only leverage the measurement data at terminal nodes to identify the state-space model of the entire pipeline network. Additionally, the dynamic process of gas cannot be expressed merely as a state transition function due to topological constraints inherent in the gas network, which must be maintained throughout the dynamic process. These characteristics present significant technical challenges in identifying the state-space model of the gas pipeline network.
III State-Space Model Identification of Gas Pipeline Network
III-A Mathematical Model of Gas Pipeline Network
To derive a more practical state-space model of a gas pipeline network, we simplify and discretize the fluid partial differential equations (1)-(2) and supplement the topology constraints of the whole network.
@@ III-A Mathematical Model of Gas Pipeline Network
where $u$ denotes the speed of sound; $A$ denotes the cross-section area of pipeline. Equation (6) holds under an approximately adiabatic condition.
Finally, based on the approximation in (3)-(5) and the replacement of variables in (6)-(7), the partial differential equations (1)-(2) are transformed into
@@
$\left(1+\frac{\lambda_{l}v_{b,l}\Delta t}{D_{l}}\right)f_{i,t}-\frac{A_{l}\Delta t}{\Delta x_{l}}\pi_{i-1,t}$ (11)
$+\left(\frac{A_{l}\Delta t}{\Delta x_{l}}-\frac{\lambda_{l}A_{l}v_{b,l}^{2}\Delta t}{2u^{2}D_{l}}\right)\pi_{i,t}=f_{i,t-1}$
where $\mathcal{L}$ and $\mathcal{S}$ denote the index set of pipelines and computation nodes, respectively; $\mathcal{S}_{l}$ denotes the index set of computation node in the pipeline $l$. For the convenience of numerical computation, each pipeline is divided into sections by several computation nodes.
@@
III-B State-Space Model of Gas Pipeline Network
@@
To demonstrate the construction of state-space model, we use a 3-node small gas system as an example, as shown in Fig. 2. Each pipeline is divided into 2 segments and there are total of 6 computation points, numbered $1^{\prime}$-$6^{\prime}$ , respectively.
@@
The dynamic function (18) of this 3-node simple gas system can be written in the following matrix form as shown in Fig. 3.
@@
The detail parameters inside the sparse matrix are defined as follows:
@@
According to the matrix dynamic function in Fig. 4, to calculate the state variables $\bm{h}_{t}$, it only needs the values in $\bm{h}_{t-1}$ that start from the second computation points. This is because the state of the first computation points of each pipeline ($1^{\prime}$ and $4^{\prime}$ in this case) are not independent. Their values are also determined by topology constraints at terminal nodes. The constant matrix $\bm{S}$ is used to select the independent states from all state variables. In this small case, the value of matrices $\bm{S}$ and $\bm{H}$ are shown in Fig. 4.
III-C Identify the State-Space Model with Physics-Informed Recurrent Network
This subsection introduces the state-space model of the pipelines using the measurement data in the dynamic process. A PIRN is proposed to identify the parameters in the state-space model (18)-(19).
@@ III-C Identify the State-Space Model with Physics-Informed Recurrent Network
Based on the state transition equation (26) and the output equation (19), the diagram of PIRN model training architecture is shown in Fig. 5. The control variables $\bm{u}_{t}$ are inputs of this system. The state variables $\bm{h}_{t}$ updates based on the input variables and the previous state $\bm{h}_{t-1}$. The outputs $\hat{\bm{y}}_{t}$ are the predicted output data and the labels $\bm{y}_{t}$ are the real measured states.
Figure 5: Diagram of PIRN framework.
@@
where $\alpha_{\theta}$ and $\alpha_{h}$ denote the learning rates of $\hat{\bm{\theta}}$ and $\bm{h}_{0}$, respectively, and they are the hyper-parameters in the model. With the PyTorch framework, we can leverage automatic differentiation to compute partial derivatives numerically, facilitating backward propagation and parameter updates in (29)-(30). However, this process involves repeated inversion of the large-scale sparse matrix $\bm{K}(\hat{\bm{\theta}})$ to obtain updated $\bm{J}$ as the parameters $\hat{\bm{\theta}}$ updated, which introduces a significant computational burden and can lead to numerical instability. To address this issue, we apply the matrix inversion lemma to update the new value of the matrix $\bm{J}$. When the parameters $\hat{\bm{\theta}}$ inside the sparse matrix $\bm{K}(\hat{\bm{\theta}})$ change to $\hat{\bm{\theta}}+\Delta\hat{\bm{\theta}}$, the matrix update can be denoted as:
@@
In summary, the proposed PIRN model exploits the memory-retention capabilities of RNNs to recover the state-space description of a gas pipeline network from sparse terminal-node measurements. Embedding the physical state-space equations within the recurrent architecture ensures interpretability and allows seamless integration into any optimal-dispatch framework. A comparison between the classic RNNs and the proposed PIRN model is summarized in TABLE I. Besides, to demonstrate the effectiveness of the identified state-space model, the optimal dispatch of the integrated power-gas system is implemented in Section IV-E. By jointly optimizing the operation of generators, GTs, GWs, P2Gs and compressors, the economic benefits of the integrated system can be fully exploited.
TABLE I: Comparison between Classic RNNs and Physics-Informed Recurrent Network
Comparison items Classic RNNs Physics-informed recurrent network
Output layer Predicted data Predict measurable states of gas
Label Ground truth Measurement data of measurable variables
Training algorithm Backward propagation through time (BPTT)
Implementation framework PyTorch/TensorFlow
III-D Training Techniques of Physics-Informed Recurrent Network
Unlike traditional neural network based AI models, the parameters in the PIRN model carry precise physical meaning. To ensure both stability and efficiency during training, several specialized techniques must be employed, including warm start, data normalization, gradient clipping and truncated backward propagation through time.
III-D1 Warm Start
A warm start avoids numerical instability during matrix inversion more efficiently than random initialization, and accelerates the training process. Since all the parameters have an explicit physical meaning, the initial values of pending parameters $\hat{\bm{\theta}}$ and initial states $\bm{h}_{0}$ can be estimated based on the prior knowledge of physical meanings. The initial value of pending parameters $\hat{\bm{\theta}}$ can be directly set as $\bm{\theta}_{0}$. The initial states $\bm{h}_{0}$ in steady state can be estimated according to the physical model of gas pipeline. According to the partial differential equations of gas pipelines, the expressions of mass flow and gas pressure distribution along the pipeline in steady state are as follows: